# Tropicalization and cluster asymptotic phenomenon of generalized Markov equations

### v1

@@ Abstract.
The generalized Markov equations are deeply connected with the generalized cluster algebras of Markov type. We construct a deformed Fock-Goncharov tropicalization for the generalized Markov equations and prove that their tropicalized tree structure is essentially the same as that of the classical Euclid tree. We then define the generalized Euclid tree and prove that it converges to the classical Euclid tree up to a scalar multiple. Moreover, by means of cluster mutations, we exhibit an asymptotic phenomenon, up to some limit $q$, between the logarithmic generalized Markov tree and the classical Euclid tree. A rationality conjecture of $q$ is then put forward. We also propose a generalized Markov uniqueness conjecture for the generalized Markov equations, which illustrates an application of the asymptotic phenomenon.
Keywords: generalized Markov equations, tropicalization, cluster mutations, asymptotic phenomenon, generalized Markov uniqueness conjecture.
@@ 1.5. Conventions
For convenience, we use the following notations.
@@ 3.2. Deformed Fock-Goncharov tropicalization
$\lim_{C\rightarrow\infty}\dfrac{\log(e^{CX_{1}},\dots,e^{CX_{n}})}{C}=\max(X_{1},\dots,X_{n}).$ (3.18)
@@ 6.2. Generalized Euclid tree arising from logarithmic generalized Markov tree
We have seen in the previous subsection that the cluster mutations on generalized Markov triples are approximately multiplications when the length of mutation sequence $\mathbf{w}$ is large enough. Then, it is natural to investigate the behavior of cluster mutations when we take logarithm of the generalized Markov tree, whose generalized Markov triples $(a,b,c)$ are replaced by $(\ln(a),\ln(b),\ln(c))$. We call such tree the logarithmic generalized Markov tree.
@@
For brevity, we denote by $\widebar{x}\coloneq\ln(x)$ for any positive real number $x$. Therefore, we have
$(\widebar{a},\widebar{b},\widebar{c})\coloneq(\ln(a),\ln(b),\ln(c)).$ (6.12)
@@
$(\widebar{a_{i+1}},\widebar{b_{i+1}},\widebar{c_{i+1}})=(\ln(a_{i+1}),\ln(b_{i+1}),\ln(c_{i+1}))=(\widebar{k_{i+1}}+\widebar{b_{i}}+\widebar{c_{i}},\widebar{b_{i}},\widebar{c_{i}}).$ (6.16)
@@
Let the conditions be same as 6.2 and denote by $\widebar{k_{\lambda}}=\ln(k_{\lambda})$. Then, the relation (6.7) is equivalent to
@@ 6.3. Main results
$q=\lim_{n\rightarrow+\infty}\frac{\ln(F_{2n-1})}{n}=\lim_{n\rightarrow+\infty}\frac{\ln(F_{2n+1})}{n+1}=2\ln(\varphi)=\ln(\frac{3+\sqrt{5}}{2}).$ (6.45)
@@ 8.2. Application: an approximate method for verification
Step $2$: For a generalized Markov triple $(a,b,c)$, without loss of generality, we may assume that $a\geq b\geq c$ and its logarithmic form is $(\ln(a),\ln(b),\ln(c))$. If the counter-example of 8.2 exists, that is $(a,b^{\prime},c^{\prime})$ with $b^{\prime}\geq c^{\prime}$ and $(b^{\prime},c^{\prime})\neq(b,c)$, then by Theorem 6.10, we have
$\ln(a)\approx q\times x_{1},$ (8.1)
where $x_{1}$ is the first component of some classical Euclid triple in $\mathcal{E}$. Then, we can fix some $n$ and there are also $3\times 2^{n-1}$ possible values of $\frac{\ln(a)}{q}$.
Step $3$: In the classical Euclid tree $\mathcal{E}$, find the possible triples $(x,y,z)$ whose first component $x$ is approximate to $\frac{\ln(a)}{q}$ and $x\geq y\geq z$. They one-to-one correspond to some unique mutation sequence $\mathbf{w}_{0}$. Finally, we can check that whether the generalized Markov triple $\mu^{\mathbf{w}_{0}}(1,1,1)$ is a counter-example.
@@
The reason why we adopt the classical Euclid tree $\mathcal{E}$ is that its structure is simpler and more transparent. Moreover, this approach significantly reduces the computational effort required for the generalized Markov equations. However, as we aim for more precise results (as $n$ increases), the complexity will correspondingly grow. In conclusion, we propose an approximate method to search for counter-examples of 8.2.

### v2

@@ Abstract.
The generalized Markov equations are deeply connected with the generalized cluster algebras of Markov type. We construct a deformed Fock-Goncharov tropicalization for the generalized Markov equations and prove that their tropicalized tree structure is essentially the same as that of the classical Euclid tree. We then define the generalized Euclid tree and prove that it converges to the classical Euclid tree up to a scalar multiple. Moreover, by means of cluster mutations, we exhibit an asymptotic phenomenon, up to some limit $q$, between the logarithmic generalized Markov tree and the classical Euclid tree. A rationality conjecture of $q$ is then put forward. We also propose a generalized uniqueness conjecture for the generalized Markov equations, which illustrates an application of the asymptotic phenomenon.
Keywords: generalized Markov equations, tropicalization, cluster mutations, asymptotic phenomenon, generalized uniqueness conjecture.
@@ 1.5. Conventions
Throughout the paper, we use the following notations.
@@ 3.2. Deformed Fock-Goncharov tropicalization
    Then, it is a semifield, which is called the tropical semifield.
@@
$\lim_{C\rightarrow\infty}\dfrac{\log(e^{CX_{1}}+\dots+e^{CX_{n}})}{C}=\max(X_{1},\dots,X_{n}).$ (3.18)
@@ 6.2. Generalized Euclid tree arising from logarithmic generalized Markov tree
We have seen in the previous subsection that the cluster mutations on generalized Markov triples are approximately multiplications when the length of mutation sequence $\mathbf{w}$ is large enough. Then, it is natural to investigate the behavior of cluster mutations when we take logarithm of the generalized Markov tree, whose generalized Markov triples $(a,b,c)$ are replaced by $(\log(a),\log(b),\log(c))$. We call such tree the logarithmic generalized Markov tree.
@@
For brevity, we denote by $\widebar{x}\coloneq\log(x)$ for any positive real number $x$. Therefore, we have
$(\widebar{a},\widebar{b},\widebar{c})\coloneq(\log(a),\log(b),\log(c)).$ (6.12)
@@
$(\widebar{a_{i+1}},\widebar{b_{i+1}},\widebar{c_{i+1}})=(\log(a_{i+1}),\log(b_{i+1}),\log(c_{i+1}))=(\widebar{k_{i+1}}+\widebar{b_{i}}+\widebar{c_{i}},\widebar{b_{i}},\widebar{c_{i}}).$ (6.16)
@@
Let the conditions be same as 6.2 and denote by $\widebar{k_{\lambda}}=\log(k_{\lambda})$. Then, the relation (6.7) is equivalent to
@@ 6.3. Main results
$q=\lim_{n\rightarrow+\infty}\frac{\log(F_{2n-1})}{n}=\lim_{n\rightarrow+\infty}\frac{\log(F_{2n+1})}{n+1}=2\log(\varphi)=\log(\frac{3+\sqrt{5}}{2}).$ (6.45)
@@ 8.2. Application: an approximate method for verification
Step $2$: For a generalized Markov triple $(a,b,c)$, without loss of generality, we may assume that $a\geq b\geq c$ and its logarithmic form is $(\log(a),\log(b),\log(c))$. If the counter-example of 8.2 exists, that is $(a,b^{\prime},c^{\prime})$ with $b^{\prime}\geq c^{\prime}$ and $(b^{\prime},c^{\prime})\neq(b,c)$, then by Theorem 6.10, we have
$\log(a)\approx q\times x_{1},$ (8.1)
where $x_{1}$ is the first component of some classical Euclid triple in $\mathcal{E}$. Then, we can fix some $n$ and there are also $3\times 2^{n-1}$ possible values of $\frac{\log(a)}{q}$.
Step $3$: In the classical Euclid tree $\mathcal{E}$, find the possible triples $(x,y,z)$ whose first component $x$ is approximate to $\frac{\log(a)}{q}$ and $x\geq y\geq z$. They one-to-one correspond to some unique mutation sequence $\mathbf{w}_{0}$. Finally, we can check that whether the generalized Markov triple $\mu^{\mathbf{w}_{0}}(1,1,1)$ is a counter-example.
@@
The reason why we adopt the classical Euclid tree $\mathcal{E}$ is that its structure is simpler and more transparent. Moreover, this approach significantly reduces the computational effort required for the generalized Markov equations. However, as we aim for more precise results (as $n$ increases), the complexity will correspondingly grow. For future investigations, it is also meaningful to provide an estimate of this complexity. In conclusion, we propose an approximate method to search for the possible counter-examples of 8.2.